# Characterizing Dynamical Stability of Stochastic Gradient Descent in Overparameterized Learning

### v2

@@ 3.1 Overview
The remainder of this paper will consist of the proofs of Theorem A and Theorem B. For the main arguments, it will be convenient to work in a local coordinate system in which $\mathcal{M}$ corresponds to a linear subspace. Such a coordinate system will be introduced in section 3.2. Theorem A (i) then follows from an elementary argument presented in section 3.3. In order to prove Theorem A (ii), in section 3.4, we employ a center-stable manifold theorem. In section 3.5, we introduce a random dynamical systems framework to treat stochastic gradient descent. Theorem B (i) is then proved in section 3.6. The proof is similar to the proof of Theorem A (i) presented in section 3.3, but the possibility for non-uniform hyperbolicity adds an additional challenge. Finally, Theorem B (ii) is proved in sections 3.7-3.9. The proof is inspired by previous work on the instability of invariant subspaces for stochastic differential equations (see e.g. Baxendale and Stroock (1988); Baxendale (1991)). While these works rely on Hörmander conditions (cf. Hörmander (1967)), we use a criterion by Le Page (1982) to establish a spectral gap for the projective semi-group. The assumptions for the criterion of Le Page are checked in section 3.7. In section 3.8, we construct a local Lyapunov function similar to the recent works by Bedrossian et al. (2022) and Blumenthal et al. (2023). The proof of Theorem B is then completed in section 3.9.
@@ 3.2 Local Coordinates
In the following, for some fixed $x^{*}\in\mathcal{M}$, we introduce a local coordinate system for a neighborhood of $x^{*}$, in which the generally non-linear manifold $\mathcal{M}$ becomes a linear subspace aligned with the coordinate axes. Using these coordinates will be helpful in all further proofs.
@@
Proof  Let $A\in\mathbb{R}^{D\times(D-N)}$ be some matrix which induces an isomorphism from $\mathbb{R}^{D-N}$ onto $\mathcal{T}(x^{*})$. Since $\mathcal{M}$ is a embedded smooth manifold, we can find neighborhoods $0\in U_{\mathcal{T}}\subseteq\mathcal{T}(x^{*})$ and $x^{*}\in\tilde{U}\subseteq\mathbb{R}^{D}$ and smooth map $\zeta:U_{\mathcal{T}}\to\mathcal{N}(x^{*})$ with $\zeta(0)=0$, $\zeta^{\prime}(0)=0$ and
@@
so in particular $\chi$ is a local diffeomorphism and (3.1) holds. Set $\hat{V}:=\chi^{-1}(\tilde{U})$ and $\hat{U}:=\chi(\hat{V})$. If we now let $\chi$ be its restriction $\chi:\hat{V}\to\hat{U}$, it is a smooth diffeomorphism, satisfying i) to iii). If $\chi$ is not bi-Lipschitz we can reassign $\hat{V}$ to a smaller neighborhood, which is pre-compact in the original $\hat{V}$. After reassigning $\hat{U}$ and restricting $\chi$ accordingly, $\chi$ will be bi-Lipschitz.
@@
The learning dynamics of (stochastic) gradient decent can, at least locally, be lifted in the new coordinates via the function $\chi$. We define functions $\hat{\mathcal{G}}_{\eta},\hat{\mathcal{G}}_{\eta,1},\dots\hat{\mathcal{G}}_{\eta,N}:V^{*}\to\hat{V}$ by
@@
Let $\tau:V^{*}\to\mathbb{N}\cup\{\infty\}$ be the maximal number $n$ for which $\hat{\mathcal{G}}^{n+1}_{\eta}(v,w)$ is well defined, i.e.
@@
where the right-hand sides are both block matrices with dimensions $((D-N)+N)\times((D-N)+N)$.
@@
Note that this lemma would hold trivially if the right hand-sides of (3.6) and (3.7) were replaced by the term $\delta\|(v,w)\|$.
Proof  It is sufficient to consider each of the functions $\hat{\mathcal{G}}_{\eta},\hat{\mathcal{G}}_{\eta,1},\dots,\hat{\mathcal{G}}_{\eta,N}$ separately, find a neighborhood $V_{\delta}$ on which the corresponding inequality from (3.6) or (3.7) holds and conclude the proof by choosing the intersection over all these $V_{\delta}$. In the following, we will only consider $\hat{\mathcal{G}}_{\eta}$ and will show how to find an $V_{\delta}$ such that (3.6) holds, as the proofs for $\hat{\mathcal{G}}_{\eta,1},\dots,\hat{\mathcal{G}}_{\eta,N}$ work analogously.
@@ 3.3 Gradient Descent - the Stable Case
In the following we will let $\Pi_{v}\in\mathbb{R}^{(D-N)\times D}$ and $\Pi_{w}\in\mathbb{R}^{(D-N)\times N}$ be the matrices which project a vector $(v,w)^{t}\in\mathbb{R}^{D}$ onto its $v$-, respectively $w$-component, i.e. in block-matrix form
@@
Figure 2: Schematic representation of the proof of Theorem A (i). The red arrows show posible sample trajectories.
@@ 3.4 Gradient Descent - the Unstable Case
In this section we will prove Theorem A (ii).
@@
In the case $\mu(x^{*})<0$ considered in the previous section, all eigenvalues of $\mathds{1}_{N}-\eta G_{x^{*}}$ lie strictly inside the unit circle. As a consequence we were able to construct an open neighborhood of $x^{*}$, such that for every initial condition in that neighborhood, GD converges to some $X_{\lim}^{\operatorname{GD}}\in\mathcal{M}$ near $x^{*}$. In the case $\mu(x^{*})>0$ we consider in this section, we only know that some eigenvalue of $\mathds{1}_{N}-\eta G_{x^{*}}$ lies strictly outside the unit circle while there could still be eigenvalues on or strictly inside the unit circle. As a consequence on should not expect to be able to construct a neighborhood $U\subseteq\mathbb{R}^{D}$ of $x^{*}$, such that GD does not converge to some $X_{\lim}^{\operatorname{GD}}\in\mathcal{M}$ near $x^{*}$ for any initial condition in $U$. Instead we will construct a neighborhood $U\subseteq\mathbb{R}^{D}$ of $x^{*}$, such that GD does not converge to some $X_{\lim}^{\operatorname{GD}}\in\mathcal{M}$ near $x^{*}$ for Lebesgue-almost any initial condition in $U$. In fact, we will show that the set of initial conditions in $U$ for which GD does converge to some $X_{\lim}^{\operatorname{GD}}\in\mathcal{M}$ near $x^{*}$ is contained in a lower dimensional manifold called the center-stable manifold [^11] While the set, we call center-stable manifold is indeed a $C^{1}$-manifold (cf. Theorem 6.2.8 in Katok and Hasselblatt (1995)), we will only show that it is the graph of a Lipschitz continuous function.. While for the preceding and all subsequent sections it is convenient to consider vectors in $\mathbb{R}^{D}$ as consisting of a $D-N$-dimensional tangential part $v$ and an $N$-dimensional transversal part $w$, here it will be more convenient to consider vectors in $\mathbb{R}^{D}$ as consisting of a center-stable part $v_{-}$ corresponding to the eigenvalues strictly inside or on the unit circle and an unstable part $v_{+}$ corresponding to the eigenvalues strictly outside the unit circle.
@@
Note that the eigenvalue 1 must appear at least with multiplicity $D-N$, corresponding to the tangential part and $\log|\bar{\mu}_{1}|=\mu(x^{*})>0$, which implies $|\bar{\mu}_{1}|>1$. Let $D_{+}>1$ be the number of eigenvalues with absolute value greater 1 and let $D_{-}=D-D_{+}$ be the number of eigenvalue with absolute value less or equal to 1, counted with multiplicity in both cases. Furthermore, let $\bar{S}\in\mathbb{R}^{D\times D}$ be an invertible matrix, such that
@@
If $\bar{\mathcal{G}}_{\eta}$ is a local diffeomorphism at 0, this result is standard (see e.g.  chapter 6.2 in Katok and Hasselblatt (1995)). In our setting, however, it is possible that $\bar{\mu}_{i}=0$ for some $i\in[D]$. Still, Theorem 3.3 follows from the arguments in Katok and Hasselblatt (1995) under mild modifications. Alternatively, stable/unstable/center manifold theorems in non-locally diffeomorphic setting have also been obtain by reducing the problem to the locally diffeomorphic case by an abstract method called inverse limits (see e.g. Ruelle and Shub (1980) or Qian et al. (2009)).
In order to prove Theorem A (ii), we require the following lemma. Since the equivalent statement will also be useful to prove instability of SGD, we formulate it to cover both GD and SGD.
@@
where $V_{+}$ and $V_{-}$ are the neighborhoods given by Theorem 3.3. Suppose $\omega\in\Omega$ is such that $X_{\lim}^{\operatorname{GD}}(\omega)\in U\cap\mathcal{M}$. Then there must exists an $m\in\mathbb{N}$, such that $\mathcal{G}_{\eta}^{n}(X_{m}^{\operatorname{GD}}(\omega))=X_{m+n}^{\operatorname{GD}(\omega)}\in U$ for all $n\in\mathbb{N}_{0}$. Equivalently, we must have $\bar{\mathcal{G}}_{\eta}^{n}(\bar{S}^{-1}\chi^{-1}(X_{m}^{\operatorname{GD}}(\omega)))\in V_{+}\times V_{-}$, for all $n\in\mathbb{N}_{0}$. By Theorem 3.3 this means that we can only have $X_{\lim}^{\operatorname{GD}}(\omega)\in U\cap\mathcal{M}$, if there exists an $m\in\mathbb{N}$ such that $X_{m}^{\operatorname{GD}}(\omega)\in\chi(\bar{S}\operatorname{graph}(\beta^{*}))$. However, since $\chi$ is a diffeomorphism and $\bar{S}$ an invertible matrix, the set $\chi(\bar{S}\operatorname{graph}(\beta^{*}))$ has Lebesgue-measure zero. By Lemma 3.4, this implies
@@ 3.5 Random Dynamical System Framework for SGD
While in sections 3.3 and 3.4 we studied the dynamics of gradient descent, the rest of the paper will be concerned with establishing the analogous results for stochastic gradient descent. For this, we will formulate SGD as a random dynamical system (see Arnold (1998) for a general introduction into the theory). Formally, for some appropriate probability space $(\Omega,\mathcal{F},\mathbb{P})$, we will introduce a map
@@
where $\theta^{n}\omega$ is the seed for the same training examples, but shifted by $n$ steps. Put more clearly, if $\omega$ generates the sequence of training examples $(\xi_{1},\xi_{2},\dots)\in[N]^{\mathbb{N}}$, then $\theta^{n}\omega$ should generate the sequence $(\xi_{n+1},\xi_{n+2},\dots)\in[N]^{\mathbb{N}}$. Since the training examples are chosen independently with identical distributions the sequences $(\xi_{1},\xi_{2},\dots)$ and $(\xi_{n+1},\xi_{n+2},\dots)$ are equal in distribution. In other words the map $\theta:\Omega\to\Omega$ leaves the probability measure $\mathbb{P}$ invariant, i.e.
@@
A map $\theta:\Omega\to\Omega$ satisfying (3.16) is called metric dynamical system. Equation (3.15) is known as the cocycle property and a map $\varphi$ satisfying it is called a cocycle over $\theta$. A pair $(\theta,\varphi)$, consisting of a metric dynamical system and a cocycle over it, is called random dynamical system (cf. Defintion 1.1.1 in Arnold (1998)).
@@
Note that, again, points $x\in\mathcal{M}$ are fixed points of $\varphi$, meaning
@@
where $\mathrm{D}_{x}\varphi_{\omega}^{(n)}(x^{*})$ denotes the Jacobian of the map $\varphi_{\omega}^{(n)}$. By the chain-rule to (3.17), using (3.18), we can compute
@@
Recall that $\chi:\hat{V}\to\hat{U}$, introduced in Lemma 3.1, defines a diffeomorphism from a neighborhood $\hat{V}$ of the origin onto a neighborhood $\hat{U}$ of $x^{*}$. Furthermore, we introduced a neighborhood $V^{*}\subseteq\hat{V}$ (cf. (3.2)), which allowed us to locally lift the maps $\mathcal{G}_{\eta,1},\dots,\mathcal{G}_{\eta,n}$ via $\chi$ to maps $\hat{\mathcal{G}}_{\eta,1},\dots,\hat{\mathcal{G}}_{\eta,n}:V^{*}\to\hat{V}$ defined by
@@
whenever the left-hand side is well defined, i.e. whenever $(n+m,\omega,v,w)\in D_{\psi}$. This turns the pair $(\theta,\psi)$ into a local random dynamical system (cf. Definition 1.2.1 in Arnold (1998)). The local cocycle $\psi$ can be seen as the local lift of $\varphi$ via $\chi$, as one can easily check that
@@
Alternatively, this can also be seen by differentiating (3.21) and using Lemma 3.1, as well as (2.17). Also, differentiating (3.20) at (0,0) yields the identity
@@
for the Lyapunov exponent. By Kingman’s sub-additive ergodic theorem [^13] Note that the fact that $\lambda(x^{*})$ does not depend on $\omega$ requires $\theta$ to be ergodic, which it is not: However, the cocycle $\Psi$ is measurable with respect to the sub-sigma algebra ^$\hat{\mathcal{F}}=\sigma(\xi_{1},\xi_{2},\dots)$. Thus we may consider the probability space $(\Omega,\hat{\mathcal{F}},\mathbb{P}|_{\hat{\mathcal{F}}})$, on which $\theta$ is ergodic, when applying the sub-additive ergodic theorem., we have

### v3

@@ 3.1 Overview
The remainder of this paper will consist of the proofs of Theorem A and Theorem B. For the main arguments, it will be convenient to work in a local coordinate system in which $\mathcal{M}$ corresponds to a linear subspace. Such a coordinate system will be introduced in Section 3.2. Theorem A (i) then follows from an elementary argument presented in Section 3.3. In order to prove Theorem A (ii), in Section 3.4, we employ a center-stable manifold theorem. In Section 3.5, we introduce a random-dynamical-system framework to treat stochastic gradient descent. Theorem B (i) is then proved in Section 3.6. The proof is similar to the proof of Theorem A (i) presented in Section 3.3, but the possibility of non-uniform hyperbolicity adds an additional challenge. Finally, Theorem B (ii) is proved in sections 3.7-3.9. The proof is inspired by previous work on the instability of invariant subspaces for stochastic differential equations (see, e.g., Baxendale and Stroock, 1988; Baxendale, 1991). While these works rely on Hörmander conditions (cf. Hörmander, 1967), we use a criterion by Le Page (1982) to establish a spectral gap for the projective semigroup. The assumptions for the criterion of Le Page are checked in Section 3.7. In Section 3.8, we construct a local Lyapunov function similar to the recent works of Bedrossian et al. (2022) and Blumenthal et al. (2023). The proof of Theorem B is then completed in Section 3.9.
@@ 3.2 Local Coordinates
In the following, for some fixed $x^{*}\in\mathcal{M}$, we introduce a local coordinate system for a neighborhood of $x^{*}$, in which the generally nonlinear manifold $\mathcal{M}$ becomes a linear subspace aligned with the coordinate axes. Using these coordinates will be helpful in all further proofs.
@@
Proof  Let $A\in\mathbb{R}^{D\times(D-N)}$ be some matrix that induces an isomorphism from $\mathbb{R}^{D-N}$ to $\mathcal{T}(x^{*})$. Since $\mathcal{M}$ is an embedded smooth manifold, we can find neighborhoods $0\in U_{\mathcal{T}}\subseteq\mathcal{T}(x^{*})$ and $x^{*}\in\tilde{U}\subseteq\mathbb{R}^{D}$ and a smooth map $\zeta:U_{\mathcal{T}}\to\mathcal{N}(x^{*})$ with $\zeta(0)=0$, $\zeta^{\prime}(0)=0$ and
@@
so in particular $\chi$ is a local diffeomorphism and (3.1) holds. Set $\hat{V}:=\chi^{-1}(\tilde{U})$ and $\hat{U}:=\chi(\hat{V})$. If we now let $\chi$ be its restriction $\chi:\hat{V}\to\hat{U}$, it is a smooth diffeomorphism, satisfying i) to iii). If $\chi$ is not bi-Lipschitz, we can reassign $\hat{V}$ to a smaller neighborhood, which is precompact in the original $\hat{V}$. After reassigning $\hat{U}$ and restricting $\chi$ accordingly, $\chi$ will be bi-Lipschitz.
@@
The learning dynamics of (stochastic) gradient decent can, at least locally, be lifted in the new coordinates via the function $\chi$. We define the functions $\hat{\mathcal{G}}_{\eta},\hat{\mathcal{G}}_{\eta,1},\dots\hat{\mathcal{G}}_{\eta,N}:V^{*}\to\hat{V}$ by
@@
Let $\tau:V^{*}\to\mathbb{N}\cup\{\infty\}$ be the maximal number $n$ for which $\hat{\mathcal{G}}^{n+1}_{\eta}(v,w)$ is well defined, that is,
@@
where the right-hand sides are both block matrices with dimensions $((D-N)+N)\times((D-N)+N)$111111i.e. the rows and the columns have been partitioned into two blocks of size $(D-N)$ and size $N$..
@@
Note that this lemma would hold trivially if the right-hand sides of (3.6) and (3.7) were replaced by the term $\delta\|(v,w)\|$.
Proof  It is sufficient to consider each of the functions $\hat{\mathcal{G}}_{\eta},\hat{\mathcal{G}}_{\eta,1},\dots,\hat{\mathcal{G}}_{\eta,N}$ separately, find a neighborhood $V_{\delta}$ in which the corresponding inequality (3.6), respectively (3.7), holds and conclude the proof by choosing the intersection over all these $V_{\delta}$. In the following, we will only consider $\hat{\mathcal{G}}_{\eta}$ and will show how to find an $V_{\delta}$ such that (3.6) holds, as the proofs for $\hat{\mathcal{G}}_{\eta,1},\dots,\hat{\mathcal{G}}_{\eta,N}$ work analogously.
@@ 3.3 Gradient Descent - the Stable Case
In the following, we will let $\Pi_{v}\in\mathbb{R}^{(D-N)\times D}$ and $\Pi_{w}\in\mathbb{R}^{(D-N)\times N}$ be the matrices that project a vector $(v,w)^{t}\in\mathbb{R}^{D}$ onto its $v$-, respectively $w$-component, i.e. in block-matrix form
@@
Figure 2: Schematic representation of the proof of Theorem A (i). The red arrows show possible sample trajectories.
@@ 3.4 Gradient Descent - the Unstable Case
In this section, we will prove Theorem A (ii).
@@
In the case $\mu(x^{*})<0$ considered in the previous section, all the eigenvalues of $\mathds{1}_{N}-\eta G_{x^{*}}$ lie strictly within the unit circle. As a consequence, we were able to construct an open neighborhood of $x^{*}$, such that for every initial condition in that neighborhood, GD converges to some $X_{\lim}^{\operatorname{GD}}\in\mathcal{M}$ near $x^{*}$. In the case $\mu(x^{*})>0$ we consider in this section, we only know that some eigenvalue of $\mathds{1}_{N}-\eta G_{x^{*}}$ lies strictly outside the unit circle while there could still be eigenvalues on or strictly inside the unit circle. As a consequence, one should not expect to be able to construct a neighborhood $U\subseteq\mathbb{R}^{D}$ of $x^{*}$, such that GD does not converge to some $X_{\lim}^{\operatorname{GD}}\in\mathcal{M}$ near $x^{*}$ for any initial condition in $U$. Instead, we will construct a neighborhood $U\subseteq\mathbb{R}^{D}$ of $x^{*}$, such that GD does not converge to some $X_{\lim}^{\operatorname{GD}}\in\mathcal{M}$ near $x^{*}$ for Lebesgue-almost any initial condition in $U$. In fact, we will show that the set of initial conditions in $U$ for which GD does converge to some $X_{\lim}^{\operatorname{GD}}\in\mathcal{M}$ near $x^{*}$ is contained in a lower dimensional manifold called the center-stable manifold [^12] While the set, we call center-stable manifold is indeed a $C^{1}$-manifold (cf. Theorem 6.2.8 in Katok and Hasselblatt, 1995), we will only show that it is the graph of a Lipschitz continuous function.. While for the preceding and all subsequent sections it is convenient to consider vectors in $\mathbb{R}^{D}$ as consisting of a $D-N$-dimensional tangential part $v$ and an $N$-dimensional transversal part $w$, here it will be more convenient to consider vectors in $\mathbb{R}^{D}$ as consisting of a center-stable part $v_{-}$ corresponding to the eigenvalues strictly inside or on the unit circle and an unstable part $v_{+}$ corresponding to the eigenvalues strictly outside the unit circle.
@@
Note that the eigenvalue 1 must appear at least with multiplicity $D-N$, corresponding to the tangential part and $\log|\bar{\mu}_{1}|=\mu(x^{*})>0$, which implies $|\bar{\mu}_{1}|>1$. Let $D_{+}>1$ be the number of eigenvalues with absolute value greater than 1 and let $D_{-}=D-D_{+}$ be the number of eigenvalues with absolute value less or equal to 1, counted with multiplicity in both cases. Furthermore, let $\bar{S}\in\mathbb{R}^{D\times D}$ be an invertible matrix, such that
@@
If $\bar{\mathcal{G}}_{\eta}$ is a local diffeomorphism at 0, this result is standard (see e.g.  chapter 6.2 in Katok and Hasselblatt, 1995). However, in our setting, it is possible that $\bar{\mu}_{i}=0$ for some $i\in[D]$. Still, Theorem 3.3 follows from the arguments in Katok and Hasselblatt (1995) with mild modifications. Alternatively, stable/unstable/center manifold theorems in non-locally diffeomorphic setting have also been obtained by reducing the problem to the locally diffeomorphic case by an abstract method called inverse limits (see, e.g., Ruelle and Shub, 1980 or Qian et al., 2009).
In order to prove Theorem A (ii), we require the following lemma. Since the equivalent statement will also be useful to prove the instability of SGD, we formulate it to cover both GD and SGD.
@@
where $V_{+}$ and $V_{-}$ are the neighborhoods given by Theorem 3.3. Suppose $\omega\in\Omega$ is such that $X_{\lim}^{\operatorname{GD}}(\omega)\in U\cap\mathcal{M}$. Then there must exist an $m\in\mathbb{N}$, such that $\mathcal{G}_{\eta}^{n}(X_{m}^{\operatorname{GD}}(\omega))=X_{m+n}^{\operatorname{GD}(\omega)}\in U$ for all $n\in\mathbb{N}_{0}$. Equivalently, we must have $\bar{\mathcal{G}}_{\eta}^{n}(\bar{S}^{-1}\chi^{-1}(X_{m}^{\operatorname{GD}}(\omega)))\in V_{+}\times V_{-}$, for all $n\in\mathbb{N}_{0}$. By Theorem 3.3 this means that we can only have $X_{\lim}^{\operatorname{GD}}(\omega)\in U\cap\mathcal{M}$, if there exists an $m\in\mathbb{N}$ such that $X_{m}^{\operatorname{GD}}(\omega)\in\chi(\bar{S}\operatorname{graph}(\beta^{*}))$. However, since $\chi$ is a diffeomorphism and $\bar{S}$ an invertible matrix, the set $\chi(\bar{S}\operatorname{graph}(\beta^{*}))$ has Lebesgue measure zero. By Lemma 3.4, this implies
@@ 3.5 Random Dynamical System Framework for SGD
While in sections 3.3 and 3.4 we studied the dynamics of gradient descent, the rest of the paper will be concerned with establishing analogous results for stochastic gradient descent. For this, we will formulate SGD as a random dynamical system (see Arnold, 1998 for a general introduction to the theory). Formally, for some appropriate probability space $(\Omega,\mathcal{F},\mathbb{P})$, we will introduce a map
@@
where $\theta^{n}\omega$ is the seed for the same training examples, but shifted by $n$ steps. Put more clearly, if $\omega$ generates the sequence of training examples $(\xi_{1},\xi_{2},\dots)\in[N]^{\mathbb{N}}$, then $\theta^{n}\omega$ should generate the sequence $(\xi_{n+1},\xi_{n+2},\dots)\in[N]^{\mathbb{N}}$. Since the training examples are chosen independently with identical distributions, the sequences $(\xi_{1},\xi_{2},\dots)$ and $(\xi_{n+1},\xi_{n+2},\dots)$ are equal in distribution. In other words, the map $\theta:\Omega\to\Omega$ leaves the probability measure $\mathbb{P}$ invariant, i.e.
@@
A map $\theta:\Omega\to\Omega$ satisfying (3.16) is called metric dynamical system. Equation (3.15) is known as the cocycle property and a map $\varphi$ satisfying it is called a cocycle over $\theta$. A pair $(\theta,\varphi)$, consisting of a metric dynamical system and a cocycle over it, is called a random dynamical system (cf. Defintion 1.1.1 in Arnold, 1998).
@@
Note that the points $x\in\mathcal{M}$ are fixed points of $\varphi$, that is,
@@
where $\mathrm{D}_{x}\varphi_{\omega}^{(n)}(x^{*})$ denotes the Jacobian of the map $\varphi_{\omega}^{(n)}$. Applying the chain rule to (3.17) and using (3.18), we get
@@
Recall that $\chi:\hat{V}\to\hat{U}$, introduced in Lemma 3.1, defines a diffeomorphism from a neighborhood $\hat{V}$ of the origin to a neighborhood $\hat{U}$ of $x^{*}$. Furthermore, we introduced a neighborhood $V^{*}\subseteq\hat{V}$ (cf. (3.2)), which allowed us to locally lift the maps $\mathcal{G}_{\eta,1},\dots,\mathcal{G}_{\eta,n}$ via $\chi$ to the maps $\hat{\mathcal{G}}_{\eta,1},\dots,\hat{\mathcal{G}}_{\eta,n}:V^{*}\to\hat{V}$ defined by
@@
whenever the left-hand side is well defined, i.e. whenever $(n+m,\omega,v,w)\in D_{\psi}$. This turns the pair $(\theta,\psi)$ into a local random dynamical system (cf. Definition 1.2.1 in Arnold, 1998). The local cocycle $\psi$ can be seen as the local lift of $\varphi$ via $\chi$, as one can easily check that
@@
Alternatively, this can also be seen by differentiating (3.21) and using Lemma 3.1, as well as (2.17). Furthermore, differentiating (3.20) at (0,0) yields the identity
@@
for the Lyapunov exponent. By Kingman’s sub-additive ergodic theorem [^14] Note that the fact that $\lambda(x^{*})$ does not depend on $\omega$ requires $\theta$ to be ergodic, which it is not. However, the cocycle $\Psi$ is measurable with respect to the sub-sigma algebra ^$\hat{\mathcal{F}}=\sigma(\xi_{1},\xi_{2},\dots)$. Thus we may consider the probability space $(\Omega,\hat{\mathcal{F}},\mathbb{P}|_{\hat{\mathcal{F}}})$, in which $\theta$ is ergodic, when applying the sub-additive ergodic theorem., we have